OpenStudy (anonymous):

Find the missing length using the pythagorean theorem. a= 15 b= 1/5a c= ?

OpenStudy (anonymous):

c^2=a^2+b^2

OpenStudy (anonymous):

How would you solve it?

OpenStudy (anonymous):

Pythagorean is straight forward; as stated above.

OpenStudy (anonymous):

What do you do with the "a" after 1/5?

OpenStudy (anonymous):

You are given value of a.

OpenStudy (anonymous):

So substitute 15 in for that a?

OpenStudy (anonymous):

Substitute 15 for the "a" to get c^2=15^2+(15/5)^2 c^2 = 225+9 = 234 take square root to find c = \[\sqrt{234}\] Simplify by taking a 9 out to make it \[3*\sqrt{26}\]. This is your answer